Suppose that the cost function for the production of a particular item is given by the funtion C(x)=2x^2-320x+12020, where x represents the number of items. How many items should be produced to minimize the cost? Explain your answer
Answer by stanbon(75887)

Suppose that the cost function for the production of a particular item is given by the funtion C(x)=2x^2-320x+12020, where x represents the number of items. How many items should be produced to minimize the cost? Explain your answer
---
For a quadratic the minimum of the function occurs where x = -b/(2a).
Ans:: x = 320/(2*2) = 80 items
---------------
Cheers,
Stan H.
------------